# S, T, U Parameters in The B-LSSM

### v1

@@ IV.1 Formalism of S T U paramaters in the B-LSSM
$Y$ $\frac{1}{6}$ $-\frac{1}{2}$ $-\frac{1}{2}$ $\frac{1}{2}$ $\frac{1}{3}$ $\frac{2}{3}$1 0 0 0
$T^{3}$ $\frac{1}{2}$ 0 0 $-\frac{1}{2}$ $\frac{1}{2}$ 0 0 $-\frac{1}{2}$ $\frac{1}{2}$ 0 0 $-\frac{1}{2}$ $\frac{1}{2}$ 0 0 $-\frac{1}{2}$ 0 0 0 0 0 0
$Y_{B}$ $\frac{1}{6}$ $-\frac{1}{2}$ 0 0 $-\frac{1}{6}$ $-\frac{1}{6}$ $\frac{1}{2}$ $\frac{1}{2}$1 1
$T^{3}\times Y$ $\frac{1}{12}$ 0 0 $-\frac{1}{12}$ $-\frac{1}{2}$ 0 0 $\frac{1}{2}$ $-\frac{1}{2}$ 0 0 $\frac{1}{2}$ $\frac{1}{2}$ 0 0 $-\frac{1}{2}$ 0 0 0 0 0 0
$T^{3}\times Y_{B}$ $\frac{1}{12}$ 0 0 $-\frac{1}{12}$ $-\frac{1}{2}$ 0 0 $\frac{1}{2}$ 0 0 0 0 0 0 0 0
$Y\times Y_{B}$ $\frac{1}{36}$ $\frac{1}{4}$ 0 0 $\frac{1}{18}$ $-\frac{1}{9}$ $-\frac{1}{2}$ 0 0 0
As can be observed in the Tab. 1, the self-energy functions, namely, the unrenormalized vacuum polarisation functions with coupling constants factored out, namely, the ($\Pi^{\prime}s$), will cancel out any divergence after traversing all the superfields. It is therefore reasonable to continue using the original definition of Peskin Peskin:1991sw :
$\alpha S\equiv$ $4e^{2}[\Pi^{\prime}_{33}(0)-\Pi^{\prime}_{3Q}(0)]$
$\alpha T\equiv$ $\frac{e^{2}}{s_{W}^{2}c_{W}^{2}\widetilde{m_{Z}}^{2}}[\Pi_{11}(0)-\Pi_{33}(0)]$ (82)
$\alpha U\equiv$ $4e^{2}[\Pi^{\prime}_{11}(0)-\Pi^{\prime}_{31}(0)]$
@@
$\Pi_{AA}=e^{2}\Pi_{QQ}$ (83a)
$\Pi_{ZA}=\frac{e^{2}c^{\prime}}{s_{W}c_{W}}(\Pi_{3Q}-s_{W}^{2}\Pi_{QQ})+es^{\prime}g_{E}\Pi_{QB}$ (83b)
$\Pi_{Z^{{}^{\prime}}A}=-\frac{e^{2}s^{\prime}}{s_{W}c_{W}}(\Pi_{3Q}-s_{W}^{2}\Pi_{QQ})+ec^{\prime}g_{E}\Pi_{QB}$ (83c)
@@
$\qquad\quad\;+2\frac{es^{\prime}c^{\prime}}{s_{W}c_{W}}g_{E}(\Pi_{3B}-s^{2}_{W}\Pi_{QB})+{s^{\prime}}^{2}g^{2}_{E}\Pi_{BB}$ (83d)
@@
$\qquad\quad\;-2\frac{es^{\prime}c^{\prime}}{s_{W}c_{W}}g_{E}(\Pi_{3B}-s^{2}_{W}\Pi_{QB})+{c^{\prime}}^{2}g^{2}_{E}\Pi_{BB}$ (83e)
@@
$\qquad\quad\;+2\frac{e({c^{\prime}}^{2}-{s^{\prime}}^{2})}{s_{W}c_{W}}g_{E}(\Pi_{3B}-s^{2}_{W}\Pi_{QB})+c^{\prime}s^{\prime}g^{2}_{E}\Pi_{BB}$ (83f)
Where $g_{E}=g_{B}+\frac{Y}{Y_{B}}g_{YB}$, $Q$ indicates the charge of photon, $T_{3}$ indicates the charge of $W_{3}$ and $B$ indicates the charge of $U(1)_{B-L}$ group. The 1PI self energy of $W^{\pm}$ is given as:
$\Pi_{WW}=\frac{e^{2}}{s_{W}^{2}}\Pi_{11}$ (84)
@@
$\Pi_{AA}=e^{2}\Pi_{QQ}$ (85a)
$\widetilde{\Pi}_{ZA}=\frac{e^{2}}{s_{W}c_{W}}(\Pi_{3Q}-s_{W}^{2}\Pi_{QQ})$ (85b)
$\widetilde{\Pi}_{ZZ}=\frac{e^{2}}{s_{W}^{2}c_{W}^{2}}(\Pi_{33}-2s_{W}^{2}\Pi_{3Q}+s_{W}^{4}\Pi_{QQ})$ (85c)
@@
$\alpha S=$ $4s_{W}c_{W}[\widetilde{\Pi}^{\prime}_{ZZ}(0)-\frac{c_{W}^{2}-s_{W}^{2}}{s_{W}c_{W}}\widetilde{\Pi}^{\prime}_{Z\gamma}(0)-\Pi^{\prime}_{\gamma\gamma}(0)]$
$\alpha T=$ $\frac{\Pi_{WW}(0)}{m^{2}_{W}}-\frac{\widetilde{\Pi}_{ZZ}(0)+\frac{2s_{W}}{c_{W}}\widetilde{\Pi}_{Z\gamma}(0)+\frac{s^{2}_{W}}{c_{W}^{2}}\Pi^{\prime}_{\gamma\gamma}(0)}{m^{2}_{Z}}$ (86)
$\alpha U=$ $4s_{W}^{2}[\Pi^{\prime}_{WW}(0)-c_{W}^{2}\widetilde{\Pi}^{\prime}_{ZZ}(0)-2s_{w}c_{W}\widetilde{\Pi}^{\prime}_{Z\gamma}(0)-s_{W}^{2}\Pi^{\prime}_{\gamma\gamma}(0)]$
@@ IV.2 The S T U paramaters in the effective-Lagrangian
$+[\frac{e}{s_{W}c_{W}}(-s_{W}^{2}Q)c^{{}^{\prime}}+g_{E}Y_{B}s^{{}^{\prime}}]\bar{\psi}\gamma^{\mu}P_{R}\psi$
@@
×$\times Z^{\mu}\bar{f}\gamma_{\mu}[T^{3}P_{L}-(s_{W}^{2}-\frac{s^{\prime}(c^{2}_{W}c^{\prime}s^{\prime}(m^{2}_{Z}-m^{2}_{Z^{\prime}})+\frac{xm_{Z}m_{Z^{\prime}}}{2})}{{c^{\prime}}^{2}m^{2}_{Z}+{s^{\prime}}^{2}m^{2}_{Z^{\prime}}})]f$ (88)
Using the effective-Lagrangian techniques given by Burgess:1993vc
$\mathcal{L}_{N,Zee}=$ $\frac{e}{\hat{s_{W}}\hat{c_{W}}}(1+\frac{\alpha T}{2})Z^{\mu}\bar{f}\gamma_{\mu}[T^{3}P_{L}-(\hat{s_{W}}^{2}+\frac{\alpha S}{4(\hat{c_{W}}^{2}-\hat{s_{W}}^{2})}-\frac{\hat{c_{W}}^{2}\hat{s_{W}}^{2}\alpha T}{\hat{c_{W}}^{2}-\hat{s_{W}}^{2}})]f$
$\mathcal{L}_{C,We\nu}=$ $-\frac{e}{\sqrt{2}\hat{s_{W}}}(1-\frac{\alpha S}{4(\hat{c_{W}}^{2}-\hat{s_{W}}^{2})}+\frac{\hat{c_{W}}^{2}\alpha T}{2(\hat{c_{W}}^{2}-\hat{s_{W}}^{2})}+\frac{\alpha U}{8\hat{s}^{2}}){W^{\pm}}^{\mu}\bar{f}\gamma_{\mu}P_{L}f$ (89)
@@
$\alpha T=$ $2[\frac{\hat{s_{W}}\hat{c_{W}}}{s_{W}c_{W}}(c^{\prime}-s^{\prime}\frac{c^{\prime}s^{\prime}(m^{2}_{Z^{\prime}}-m^{2}_{Z})}{{c^{\prime}}^{2}m_{Z}^{2}+{s^{\prime}}^{2}m_{Z^{\prime}}^{2}})-1]$
$\alpha S=$ $4\hat{c_{W}}^{2}\hat{s_{W}}^{2}\alpha T+4(\hat{c_{W}}^{2}-\hat{s_{W}}^{2})(s_{W}^{2}-\hat{s_{W}}^{2}-\frac{s^{\prime}}{c^{\prime}}\frac{c_{W}^{2}c^{\prime}s^{\prime}(m^{2}_{Z^{\prime}}-m^{2}_{Z})+\frac{xm_{Z}m_{Z^{\prime}}}{2}}{{c^{\prime}}^{2}m^{2}_{Z}+{s^{\prime}}^{2}m_{Z^{\prime}}^{2}})$
$\alpha U=$ $8\hat{s_{W}}^{2}(\frac{\hat{s_{W}}}{s_{W}}-1+\frac{\alpha S}{4(\hat{c_{W}}^{2}-\hat{s_{W}}^{2})}-\frac{\hat{c_{W}}^{2}\alpha T}{2(\hat{c_{W}}^{2}-\hat{s_{W}}^{2})})$ (91)
@@
$\qquad\!\!=1-\hat{s_{W}}^{2}+2\hat{s_{W}}\Delta s=\hat{c_{W}}^{2}+2\hat{s_{W}}\Delta s$ (92)
@@
${s^{\prime}}^{2}(\frac{m^{2}_{Z^{\prime}}}{m^{2}_{Z}}-1)=-2\frac{\hat{s_{W}}}{\hat{c_{W}}^{2}}\Delta s$ (93)
Combining Eq. (91) and Eq. (93), one can obtain the following results through small-scale analysis:
$\alpha T=$ $2\frac{\Delta s}{\hat{s_{W}}}-{s^{\prime}}^{2}$ (94)
$\alpha S+\alpha U=$ $8\hat{s_{W}}\Delta s-4s^{\prime}\frac{\hat{c_{W}}\hat{s_{W}}}{e}g_{B}$ (95)
$2\hat{s_{W}}^{2}\alpha S+(\hat{s_{W}}^{2}-\hat{c_{W}}^{2})\alpha U=$ $4\hat{c_{W}}^{2}\hat{s_{W}}^{2}\alpha T+8\hat{s_{W}}(\hat{s_{W}}^{2}-\hat{c_{W}}^{2})\Delta s$ (96)
@@ IV.3 Gauge-invariant electroweak interaction parameters
$\alpha(S)_{GI}=$ $\alpha(S)_{C}+8g^{2}_{2}s_{W}c_{W}I^{\prime}_{WW}(0)({c^{\prime}}^{2}m_{Z}^{2}+{s^{\prime}}^{2}m_{Z^{\prime}}^{2})$ (97)
$\alpha(T)_{GI}=$ $\alpha(T)_{C}+4g^{2}_{2}[s_{W}^{2}I_{W\gamma}(0)+c_{W}^{2}({c^{\prime}}^{2}I_{WZ}(0)+{s^{\prime}}^{2}I_{WZ^{\prime}}(0))-I_{WW}(0)]$ (98)
@@
$-I^{\prime}_{WW}(0)]-[s_{W}^{2}I_{W\gamma}(0)+c_{W}^{2}({c^{\prime}}^{2}I_{WZ}(0)+{s^{\prime}}^{2}I_{WZ^{\prime}}(0))-I_{WW}(0)]\}$ (99)
Formally speaking, our results are extensions of Degrassi’s results Degrassi:1993kn based on the SM. The divergence of both has been cancel out, This proves the viewpoint of the first subsection of this section.
V The divergence cancellation of S T U parameters in the Neutralino-Chargino sector and srquark sector
In the context of one-loop diagram calculations, particularly in the context of shell renormalization, the elimination of divergences serves as a crucial indicator for the verification of calculation results. This chapter presents the divergence cancellation of $S$, $T$ and $U$ parameters in the neutralino-chargino sector and squark sector, with the objective of further validating the theoretical framework.

### v2

@@ IV.1 Formalism of S T U paramaters in the B-LSSM
$Y$ $\frac{1}{6}$ $\frac{1}{2}$ $\frac{1}{2}$ $\frac{1}{2}$ $\frac{1}{3}$ $\frac{2}{3}$ 1-1 0 0 0
$T^{3}$ $\left(\begin{smallmatrix}\frac{1}{2}&0\\ 0&-\frac{1}{2}\end{smallmatrix}\right)$ $\left(\begin{smallmatrix}\frac{1}{2}&0\\ 0&-\frac{1}{2}\end{smallmatrix}\right)$ $\left(\begin{smallmatrix}\frac{1}{2}&0\\ 0&-\frac{1}{2}\end{smallmatrix}\right)$ $\left(\begin{smallmatrix}\frac{1}{2}&0\\ 0&-\frac{1}{2}\end{smallmatrix}\right)$ 0 0 0 0 0 0
$Y_{B}$ $\frac{1}{6}$ $\frac{1}{2}$ 0 0 $\frac{1}{6}$ $\frac{1}{6}$ $\frac{1}{2}$ $\frac{1}{2}$ 1-1 11
$T^{3}\times Y$ $\left(\begin{smallmatrix}\frac{1}{12}&0\\ 0&-\frac{1}{12}\end{smallmatrix}\right)$ $\left(\begin{smallmatrix}-\frac{1}{2}&0\\ 0&\frac{1}{2}\end{smallmatrix}\right)$ $\left(\begin{smallmatrix}-\frac{1}{2}&0\\ 0&\frac{1}{2}\end{smallmatrix}\right)$ $\left(\begin{smallmatrix}\frac{1}{2}&0\\ 0&-\frac{1}{2}\end{smallmatrix}\right)$ 0 0 0 0 0 0
$T^{3}\times Y_{B}$ $\left(\begin{smallmatrix}\frac{1}{12}&0\\ 0&-\frac{1}{12}\end{smallmatrix}\right)$ $\left(\begin{smallmatrix}-\frac{1}{2}&0\\ 0&\frac{1}{2}\end{smallmatrix}\right)$ 0 0 0 0 0 0 0 0
$Y\times Y_{B}$ $\frac{1}{36}$ $\frac{1}{4}$ 0 0 $\frac{1}{18}$ $\frac{1}{9}$ $\frac{1}{2}$ 0 0 0
As can be observed in the Tab. 1, the self-energy functions, namely, the unrenormalized vacuum polarisation functions with coupling constants factored out, namely, the ($\Pi^{\prime}s$), will cancel out any divergence after traversing all the superfields. It is therefore reasonable to continue using the original definition of Peskin and TakeuchiPeskin:1991sw :
$\alpha S\equiv$ $4e^{2}[\Pi^{\prime}_{33}(0)-\Pi^{\prime}_{3Q}(0)],$
$\alpha T\equiv$ $\frac{e^{2}}{s_{W}^{2}c_{W}^{2}\widetilde{m_{Z}}^{2}}[\Pi_{11}(0)-\Pi_{33}(0)],$ (31)
$\alpha U\equiv$ $4e^{2}[\Pi^{\prime}_{11}(0)-\Pi^{\prime}_{31}(0)],$
@@
$\Pi_{AA}=e^{2}\Pi_{QQ}.$ (32a)
$\Pi_{ZA}=\frac{e^{2}c^{\prime}}{s_{W}c_{W}}(\Pi_{3Q}-s_{W}^{2}\Pi_{QQ})+es^{\prime}g_{E}\Pi_{QB}.$ (32b)
$\Pi_{Z^{{}^{\prime}}A}=-\frac{e^{2}s^{\prime}}{s_{W}c_{W}}(\Pi_{3Q}-s_{W}^{2}\Pi_{QQ})+ec^{\prime}g_{E}\Pi_{QB}.$ (32c)
@@
$\qquad\quad\;+2\frac{es^{\prime}c^{\prime}}{s_{W}c_{W}}g_{E}(\Pi_{3B}-s^{2}_{W}\Pi_{QB})+{s^{\prime}}^{2}g^{2}_{E}\Pi_{BB}.$ (32d)
@@
$\qquad\quad\;-2\frac{es^{\prime}c^{\prime}}{s_{W}c_{W}}g_{E}(\Pi_{3B}-s^{2}_{W}\Pi_{QB})+{c^{\prime}}^{2}g^{2}_{E}\Pi_{BB}.$ (32e)
@@
$\qquad\quad\;+2\frac{e({c^{\prime}}^{2}-{s^{\prime}}^{2})}{s_{W}c_{W}}g_{E}(\Pi_{3B}-s^{2}_{W}\Pi_{QB})+c^{\prime}s^{\prime}g^{2}_{E}\Pi_{BB}.$ (32f)
Where $g_{E}=g_{B}+\frac{Y}{Y_{B}}g_{YB}$, $Q$ indicates the electric charge, $T_{3}$ indicates the charge of $W_{3}$ and $B$ indicates the charge of $U(1)_{B-L}$ group. The 1PI self energy of $W^{\pm}$ is given as:
$\Pi_{WW}=\frac{e^{2}}{s_{W}^{2}}\Pi_{11}.$ (33)
@@
$\Pi_{AA}=e^{2}\Pi_{QQ}.$ (34a)
$\widetilde{\Pi}_{ZA}=\frac{e^{2}}{s_{W}c_{W}}(\Pi_{3Q}-s_{W}^{2}\Pi_{QQ}).$ (34b)
$\widetilde{\Pi}_{ZZ}=\frac{e^{2}}{s_{W}^{2}c_{W}^{2}}(\Pi_{33}-2s_{W}^{2}\Pi_{3Q}+s_{W}^{4}\Pi_{QQ}).$ (34c)
@@
$\alpha S=$ $4s_{W}c_{W}[\widetilde{\Pi}^{\prime}_{ZZ}(0)-\frac{c_{W}^{2}-s_{W}^{2}}{s_{W}c_{W}}\widetilde{\Pi}^{\prime}_{Z\gamma}(0)-\Pi^{\prime}_{\gamma\gamma}(0)],$
$\alpha T=$ $\frac{\Pi_{WW}(0)}{m^{2}_{W}}-\frac{\widetilde{\Pi}_{ZZ}(0)+\frac{2s_{W}}{c_{W}}\widetilde{\Pi}_{Z\gamma}(0)+\frac{s^{2}_{W}}{c_{W}^{2}}\Pi^{\prime}_{\gamma\gamma}(0)}{m^{2}_{Z}},$ (35)
$\alpha U=$ $4s_{W}^{2}[\Pi^{\prime}_{WW}(0)-c_{W}^{2}\widetilde{\Pi}^{\prime}_{ZZ}(0)-2s_{w}c_{W}\widetilde{\Pi}^{\prime}_{Z\gamma}(0)-s_{W}^{2}\Pi^{\prime}_{\gamma\gamma}(0)].$
@@ IV.2 The S T U paramaters in the effective-Lagrangian
$+[\frac{e}{s_{W}c_{W}}(-s_{W}^{2}Q)c^{{}^{\prime}}+g_{E}Y_{B}s^{{}^{\prime}}]\bar{\psi}\gamma^{\mu}P_{R}\psi.$
@@
$\times Z^{\mu}\bar{f}\gamma_{\mu}[T^{3}P_{L}-(s_{W}^{2}-\frac{s^{\prime}(c^{2}_{W}c^{\prime}s^{\prime}(m^{2}_{Z}-m^{2}_{Z^{\prime}})+\frac{xm_{Z}m_{Z^{\prime}}}{2})}{{c^{\prime}}^{2}m^{2}_{Z}+{s^{\prime}}^{2}m^{2}_{Z^{\prime}}})]f.$ (37)
Current experimental constraints on the $Z^{\prime}$ boson impose a lower bound on the ratio $M_{Z^{\prime}}/g_{B}>6~\text{TeV}$ Cacciapaglia:2006pk ; Carena:2004xs . Theoretical analyses indicate that direct contributions from $Z^{\prime}$-mediated neutral currents to electroweak-scale observables are suppressed to a precision level below $\mathcal{O}(10^{-3})$, as determined by the interplay of the $Z^{\prime}$-boson mass scale and its coupling strength. Consequently, these neutral current effects can be safely neglected in calculations of electroweak precision parameters within the current experimental resolution. This suppression arises dominantly from the decoupling behavior of the heavy $Z^{\prime}$, whose contributions to low-energy processes scale as $\sim(M_{Z}^{2}/M_{Z^{\prime}}^{2})\times g_{B}^{2}$. The stringent experimental bound on $M_{Z^{\prime}}/g_{B}$ ensures that such terms remain subleading compared to SM electroweak radiative corrections. Using the effective-Lagrangian techniques given by Burgess:1993vc
$\mathcal{L}_{N,Zee}=$ $\frac{e}{\hat{s_{W}}\hat{c_{W}}}(1+\frac{\alpha T}{2})Z^{\mu}\bar{f}\gamma_{\mu}[T^{3}P_{L}-(\hat{s_{W}}^{2}+\frac{\alpha S}{4(\hat{c_{W}}^{2}-\hat{s_{W}}^{2})}-\frac{\hat{c_{W}}^{2}\hat{s_{W}}^{2}\alpha T}{\hat{c_{W}}^{2}-\hat{s_{W}}^{2}})]f,$
$\mathcal{L}_{C,We\nu}=$ $-\frac{e}{\sqrt{2}\hat{s_{W}}}(1-\frac{\alpha S}{4(\hat{c_{W}}^{2}-\hat{s_{W}}^{2})}+\frac{\hat{c_{W}}^{2}\alpha T}{2(\hat{c_{W}}^{2}-\hat{s_{W}}^{2})}+\frac{\alpha U}{8\hat{s}^{2}}){W^{\pm}}^{\mu}\bar{f}\gamma_{\mu}P_{L}f.$ (38)
@@
$\alpha T=$ $2[\frac{\hat{s_{W}}\hat{c_{W}}}{s_{W}c_{W}}(c^{\prime}-s^{\prime}\frac{c^{\prime}s^{\prime}(m^{2}_{Z^{\prime}}-m^{2}_{Z})}{{c^{\prime}}^{2}m_{Z}^{2}+{s^{\prime}}^{2}m_{Z^{\prime}}^{2}})-1],$
$\alpha S=$ $4\hat{c_{W}}^{2}\hat{s_{W}}^{2}\alpha T+4(\hat{c_{W}}^{2}-\hat{s_{W}}^{2})(s_{W}^{2}-\hat{s_{W}}^{2}-\frac{s^{\prime}}{c^{\prime}}\frac{c_{W}^{2}c^{\prime}s^{\prime}(m^{2}_{Z^{\prime}}-m^{2}_{Z})+\frac{xm_{Z}m_{Z^{\prime}}}{2}}{{c^{\prime}}^{2}m^{2}_{Z}+{s^{\prime}}^{2}m_{Z^{\prime}}^{2}}),$
$\alpha U=$ $8\hat{s_{W}}^{2}(\frac{\hat{s_{W}}}{s_{W}}-1+\frac{\alpha S}{4(\hat{c_{W}}^{2}-\hat{s_{W}}^{2})}-\frac{\hat{c_{W}}^{2}\alpha T}{2(\hat{c_{W}}^{2}-\hat{s_{W}}^{2})}).$ (40)
@@
$\qquad\!\!=1-\hat{s_{W}}^{2}+2\hat{s_{W}}\Delta s=\hat{c_{W}}^{2}+2\hat{s_{W}}\Delta s.$ (41)
@@
${s^{\prime}}^{2}(\frac{m^{2}_{Z^{\prime}}}{m^{2}_{Z}}-1)=-2\frac{\hat{s_{W}}}{\hat{c_{W}}^{2}}\Delta s.$ (42)
Combining Eq. (40) with Eq. (42), the following results can be derived through a perturbative analysis:
$\alpha T=$ $2\frac{\Delta s}{\hat{s_{W}}}-{s^{\prime}}^{2}.$ (43)
$\alpha S+\alpha U=$ $8\hat{s_{W}}\Delta s-4s^{\prime}\frac{\hat{c_{W}}\hat{s_{W}}}{e}g_{B}.$ (44)
$2\hat{s_{W}}^{2}\alpha S+(\hat{s_{W}}^{2}-\hat{c_{W}}^{2})\alpha U=$ $4\hat{c_{W}}^{2}\hat{s_{W}}^{2}\alpha T+8\hat{s_{W}}(\hat{s_{W}}^{2}-\hat{c_{W}}^{2})\Delta s.$ (45)
@@ IV.3 Gauge-invariant electroweak interaction parameters
$\alpha(S)_{GI}=$ $\alpha(S)_{C}+8g^{2}_{2}s_{W}c_{W}I^{\prime}_{WW}(0)({c^{\prime}}^{2}m_{Z}^{2}+{s^{\prime}}^{2}m_{Z^{\prime}}^{2}).$ (46)
$\alpha(T)_{GI}=$ $\alpha(T)_{C}+4g^{2}_{2}[s_{W}^{2}I_{W\gamma}(0)+c_{W}^{2}({c^{\prime}}^{2}I_{WZ}(0)+{s^{\prime}}^{2}I_{WZ^{\prime}}(0))-I_{WW}(0)].$ (47)
@@
$-I^{\prime}_{WW}(0)]-[s_{W}^{2}I_{W\gamma}(0)+c_{W}^{2}({c^{\prime}}^{2}I_{WZ}(0)+{s^{\prime}}^{2}I_{WZ^{\prime}}(0))-I_{WW}(0)]\}.$ (48)
The subscript C here follows the definition in the literature Degrassi:1993kn , the $S$, $T$ and $U$ parameters exclude the pinch terms that ensure gauge independence.Formally speaking, our results are extensions of Degrassi’s results Degrassi:1993kn based on the SM. The divergences in both expressions cancel out. This confirms the assertion made in the first subsection of this section.
V The divergence cancellation of S T U parameters in the Neutralino-Chargino sector and squark sector
In the context of one-loop diagram calculations, particularly in the context of renormalization, the elimination of divergences serves as a crucial indicator for the verification of calculation results. This section presents the divergence cancellation of the S, T, and U parameters in the neutralino-chargino and squark sectors, with the objective of further validating the theoretical framework.